M​ is​ the​ mid​point​ of​ AC, find​ the​ coordinates​ of​ the​ missing​ end​point


1​ A​(-9,6) and​ M(5, 1)

2​ C​(1,-8) and​ M​ (-8,-4)


1
Expert's answer
2021-10-12T13:59:00-0400

Midpoint formula:

"M(x,y) = \frac{x_1+x_2}{2}, \frac{y_1+y_2}{2}"

1.

"(5,1) = \frac{-9+x_2}{2}, \frac{6+y_2}{2} \\

5 = \frac{-9+x_2}{2} \\

10 = -9 + x_2 \\

x_2 = 19 \\

1 = \frac{6+y_2}{2} \\

2 = 6+ y_2 \\

y_2 = -4

C =(19,-4)"

2.

"(-8,-4) = \frac{x_1+1}{2}, \frac{y_1-8}{2} \\

-8 = \frac{x_1+1}{2} \\

-16 = x_1 + 1 \\

x_1 = -17 \\

-4 = \frac{y_1-8}{2} \\

-8 = y_1 -8 \\

y_1=0 \\

A = (-17,0)"